A 0.835 sample of hydrated calcium nitrate was heated carefully it has changed completely to the anhydrous form. After heating, the sample mass was recorded as 0.580g. Determine the value of x(must be a whole number)


1
Expert's answer
2021-05-11T05:57:33-0400

Solution:

anhydrous form of calcium nitrate = Ca(NO3)2 (0.580 g)

hydrated calcium nitrate = Ca(NO3)2 ∙ xH2O (0.835 g)

Therefore,

Mass of xH2O = 0.835 g - 0.580 g = 0.255 g


The molar mass of Ca(NO3)2 is 164.088 g/mol

Therefore,

Moles of Ca(NO3)2 = [0.580 g Ca(NO3)2] × [1 mol Ca(NO3)2 / 164.088 g Ca(NO3)2] = 0.00353 mol

Moles of Ca(NO3)2 = 0.00353 mol

The molar mass of H2O is 18.0153 g/mol

Therefore,

Moles of H2O = [0.255 g H2O] × [1 mol H2O / 18.0153 g H2O] = 0.01415 mol

Moles of H2O = 0.01415 mol


Ca(NO3)2: 0.00353 / 0.00353 = 1

H2O: 0.01415 / 0.00353 = 4

Thus,

The value of x is 4.

Ca(NO3)2 ∙ xH2O = Ca(NO3)2 ∙ 4H2O


Answer: The value of x is 4